OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

\[-2(x+3)=-2(x+1)-4\] \[-2x-6=-2x-2-4\] \[-2x+2x=-2-4+6\] \[0x=0\] \[x \in(-\infty;+\infty)\]

OpenStudy (anonymous):

what does it = ? ... @Gretacig

OpenStudy (anonymous):

it means that you can take any x, it will be fit, because there is multiplication of 0.

OpenStudy (anonymous):

so it would be all real numbers ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks so much! (: